OpenStudy (anonymous):

what is the best approximation of the projection of (2,6) onto (5,-1)? A. 0.10(2,6) B. 0.15(5,-1) C. 0.10(5,-1) D. 0.15(2,6)

OpenStudy (aum):

Projection of vector A onto B is given by the formula: \[ \frac{A \cdot B}{|B|} * \frac{B}{|B|} \\ \text{ } \\ \text{Note: }\frac{A \cdot B}{|B|} \text{ is the scalar projection of A onto B or the component of A }\\ \text{in the direction of B and } \frac{B}{|B|} ~\text{is the unit vector in the direction of B. } \\ \text{Here, }A = <2, 6>; ~~B = <5, -1> \\ \text{Plug A and B into the formula, do the dot product and the rest and simplify.} \]